OpenStudy (lilmike234):

What is the center and radius of the circle (x+4)^2+(y-2)^2=16

jimthompson5910 (jim_thompson5910):

Hint: Rewrite \[\Large (x+4)^2+(y-2)^2=16\] as \[\Large (x-(-4))^2+(y-2)^2=4^2\] and compare that to the general form\[\Large (x-h)^2+(y-k)^2=r^2\]

OpenStudy (lilmike234):

So the center is (4,-2) and the radius is 16

OpenStudy (usukidoll):

for the r what is the square root of 16?

OpenStudy (lilmike234):

4

OpenStudy (usukidoll):

yes... so our radius is actually 4. our center is in the form of (h,k). I think opposites signs are needed. I have to recheck this one fast one.

OpenStudy (lilmike234):

So the center is (-4,2)

OpenStudy (usukidoll):

ah it's like solving for x and solving for y yeah(-4,2) is the center... radius is 4

jimthompson5910 (jim_thompson5910):

\[\Large (x-({\color{red}{-4}}))^2+(y-{\color{blue}{2}})^2={\color{green}{4}}^2\] \[\Large (x-{\color{red}{h}})^2+(y-{\color{blue}{k}})^2={\color{green}{r}}^2\]

OpenStudy (usukidoll):

ok... it's (-4,2) for center radius is 4 ... whew just double checking. I"m 75% sleepy